OpenStudy (anonymous):

The difference between a number and eleven is three more than half the number. what is it the algebraic equation ???????

OpenStudy (radar):

Use your algebra training. Let x = the number then, x-11 = 3 + x/2 would be the difference between the number and eleven is (=) three more than 1/2 the number.

OpenStudy (radar):

The question did not ask for the value of the number but: x -11 = 3 + x/2 subtract x/2 from both sides getting: x - x/2 -11 = 3 Now add 11 to both sides getting x - x/2 = 14 Multiply thru by 2 getting 2x - x = 28 Combine like terms x = 28 the solution for x Proof: 28 - 11 = 3 + 28/2 17 = 17